A card is selected at random from a standard deck. What are the odds that the card is not the ace of spades?
Answer by dabanfield(803)

The sum of the probabilities of an event happening and not happening is 1. That is P(e) + P(not e) = 1.
The probability of drawing an ace of spades is 1/52 (one card from a deck of 52).
So the probability of not drawing the ace of spades is 1-(1/52) = 51/52.